OpenStudy (anonymous):

write the expression as a single logarithm. 4 log x – 6 log (x + 2) (1 point) A. 24 log B. log x4(x + 2)6 C. log x (x + 2)24 D. none of these

OpenStudy (mathmate):

Using the laws of log, a log(x)=log(x^a) and log(a)-log(b)=log(a/b) Do you find an answer that represents the given expression?

OpenStudy (anonymous):

first of all use a*log(x)=log(x^a) for : 4 log x (I) and 6 log (x + 2) (II) then use log(a)-log(b)=log(a/b) for equation (I) and (II)

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

do you solve it?